# Further Characterization of the JadePix-3 CMOS pixel sensor for the CEPC vertex detector

### v2

@@ Abstract
The Circular Electron-Positron Collider (CEPC), a proposed next-generation $e^{+}e^{-}$ collider to enable high-precision studies of the Higgs boson and potential new physics, imposes rigorous demands on detector technologies, particularly the vertex detector. JadePix-3 is a prototype Monolithic Active Pixel Sensor (MAPS) designed for the CEPC vertex detector. This paper presents a detailed laboratory-based characterization of the JadePix-3 sensor, focusing on the previously under-explored effects of substrate reverse bias voltage on key performance metrics: charge collection efficiency, average cluster size, and detection efficiency. Systematic testing demonstrated that JadePix-3 operates reliably under reverse bias, exhibiting a reduced input capacitance, an expanded depletion region, enhanced charge collection efficiency, and a lower fake-hit rate. These findings confirm the sensor’s potential for high-precision particle tracking and vertexing at the CEPC while offering valuable references for future iterational R&D of the JadePix series.
@@ 1.1 Overview of JadePix-3 design
The JadePix-3 sensor, developed with TowerJazz 180 nm CMOS imaging technology, covers a pixel matrix consisting of 512 rows and 192 columns. The sensor is divided into four distinguishable sectors in the column direction. While the row pitch is uniformly set at 16 $\rm{\upmu m}$ across all four sectors, the column pitch varies between 23.11 $\rm{\upmu m}$ in sector 2 and 26 $\rm{\upmu m}$ in sectors 0, 1, and 3. The variation in the column pitch arises from the implementation of different digital logic blocks within the pixels. These blocks were designed to evaluate the performance of R-S flipflops versus D flipflops, as well as the inclusion or exclusion of a MASK configuration register, requiring adjustments to the pixel layout. All these differences are summarized in Tab. 2.
@@
Table 2: Analog front-ends and digital logic blocks for different sectors.
Sector pixel layout digital logic block Remarks
0 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 R-S flipflop
1 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V1 D flipflop, w/o MASK
2 16 $\times$ 23.11 $\rm{\upmu m^{2}}$ DGT_V2 R-S flipflop, w/o MASK
3 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 Increased current for the front-end
@@ 1.2 Characterization Motivation
Figure 2: Circuit topology for the JadePix-3 analog front-end [5].
The performance of MAPS is significantly influenced by the bias voltage applied to the substrate. This bias has a multifaceted impact: on one hand, it directly increases the size of the depletion zone and reduces the sensor’s input capacitance [7]. On the other hand, it also affects the characteristics of the integrated circuitry; specifically, as the substrate bias becomes more pronounced, the threshold of NMOS transistors within the pixel circuitry (shown in Fig. 2) tends to increase due to the body effect, consequently altering the circuit’s operating status. Therefore, adjusting the circuit’s operating point after applying substrate bias and investigating the resulting impact on these key performance metrics becomes essential. These metrics are vital for the R&D of the CEPC vertex detector.
@@ 2.2 Working-point Selection
The operational parameters of the pixels—threshold, fixed pattern noise (FPN), and temporal noise (TN)—were characterized using the standard “S-curve” method [8]. This technique involves injecting a series of electrical pulses of increasing charge ($Q_{\rm injected}$) into each pixel and measuring the hit response. The injected charge is regulated by an external voltage pulse, $V_{\rm pulse}$, according to the relation $Q_{\rm injected}=V_{\rm pulse}\cdot C_{\rm test}$, where $C_{\rm test}$ is the injection capacitor with a designed value of 0.18 fF. The resulting S-shaped curve for each pixel is then analyzed to extract its specific threshold and noise characteristics. The final threshold value is converted to electrons via the value of $C_{\rm test}$.
@@ 0 V 400 500 1307 722
The threshold distributions for these target working-points are presented in Fig. 4. It is observed that these distributions at non-zero reverse biases are noticeably broader compared to zero bias, indicating increased pixel-to-pixel threshold variations. This broadening of the threshold distributions is quantified in Fig. 5, which plots the FPN as a function of bias voltage. The FPN is observed to increase with reverse bias, reflecting the growing pixel-to-pixel variations. The corresponding TN analysis is shown in Fig. 5. An increase in TN is observed under reverse bias conditions. The underlying reasons for this behavior will be discussed in Section 3.2 along with the change of the OUT_A signal amplitude.
@@ 3.2 Input Capacitance
The two methods provide complementary insights. While Method 1 directly captures the amplitude response at a fixed pulse voltage, Method 2, by contrast, keeps the amplitude-dependent $f_{\rm gain}$ factor via maintaining the amplitude value constant, which offers a more reliable $C_{\rm effective}$ result with increasing reverse bias voltage.
Results for these two methods after calculation are shown in Fig. 8. In both methods, the input capacitance can be reduced by more than 50% when going from $V_{\rm sub}=0~{\rm V}$ to $V_{\rm sub}=-2~{\rm V}$, corresponding to the rapid expansion of the depletion region. This reduction is very important as smaller input capacitance increases the charge-voltage conversion coefficient at the input of the front-end circuit for a given charge. As a consequence of the increased charge-voltage conversion, the amplitude of the analog signal (OUT_A) increases significantly with reverse bias. In Section 2.2.1, to keep the input-referred threshold constant, the VCASN was tuned at different reverse bias voltage. The tuning of VCASN actually lowers the effective OUT_A baseline which is relative to the voltage required to switch the transistor M9 on. The increase of FPN and TN in Fig. 5 5, respectively, may be related to the tuning of VCASN because the same effect observed in a dedicated VCASN scan shown in Ref. [5] (Fig. 13).
@@ 3.3 Fake-hit Rate Test
JadePix-3 employs both rolling shutter (RS) and global shutter (GS) readout sequence modes. In the conventional rolling shutter mode, one row is activated for reading at a time [11]. This process requires 192 ns to transmit hit data and reset the pixels, and scanning the entire 512-row matrix takes 98.3 $\rm{\upmu s}$. While in global shutter, all 512 rows are frozen for readout after a period of acquisition time. To compare the two readout sequence modes, the acquisition time of GS mode is set equivalent to 98.3 $\rm{\upmu s}$ and the fake-hit rate for each sector $i$ in both modes can be calculated using a unified formula:
@@
where $N_{i}^{\text{mode}}$ represents the total number of false hits in sector $i$, $N_{\text{frames}}^{\text{mode}}$ is the number of frames collected, and $N_{\text{pixels}}=512\times 48$ is the number of pixels per sector. For RS mode, the test system can readout continuously while for the GS mode the test system has to launch one readout for one frame, which limits the total number of frames that can be collected in the same amount of time. In the absence of external sources, a total of 100,000 frames ($N_{\text{frames}}^{\text{RS}}=100,000$) are collected in the RS mode, while 1,000 frames ($N_{\text{frames}}^{\text{GS}}=1,000$) are collected in the GS mode.
The sensitivity limit is defined as the lowest measurable fake-hit in the measurement and is determined by the number of frames for these two modes. Based on Eq. 3, this results in a sensitivity limit of $\rm 4\times 10^{-10}$ for RS mode and $\rm 4\times 10^{-8}$ for GS mode.
@@ 3.4 Radioactive Source Test
Two radioactive sources, $^{90}\rm Sr$ and $^{55}\rm Fe$, were employed to assess the sensor’s response to external ionizing radiation. $^{90}\rm Sr$, a $\beta$ emitter, generates electrons that traverse the sensor and deposit charge over an extended region. This distributed ionization is particularly sensitive to changes in the sensor’s depletion characteristics and charge-sharing behavior under reverse bias, thereby providing insights into the sensor’s charge collection efficiency. $^{55}\rm Fe$ emits 5.9 keV X-rays that interact with silicon primarily via the photoelectric effect and deposit an ionization charge of 1640 $e^{-}$ in a small point-like spatial region compared to the dimensions of the sensor, resulting in highly localized charge deposition. This characteristic makes $^{55}\rm Fe$ an ideal probe for evaluating the sensor’s response to concentrated ionization, providing a signal whose magnitude is comparable to that of a Minimum Ionizing Particle (MIP) [12].
@@ 3.5 Infrared Laser Test
Following the configuration of the laser test platform outlined in Section 2.1, a systematic investigation of the in-pixel detection efficiency was performed using an infrared laser beam incident from the backside of the chip. The laser operates at a wavelength of 1064 nm. The laser’s pulse energy is adjustable between 0 and 60 pJ, modulated via an attenuation factor ranging from 0% tune (maximum pulse energy) to 100% tune (minimum pulse energy), controlled through the Laser Power Tune function of the system. For this laser measurement, the selected pixel is located in sector 2, which has the smallest pixel pitch among the sectors (see Tab. 2).
@@ 3.5.2 In-Pixel Detection Efficiency
Previous beam test results [6] indicated a non-uniform in-pixel detection efficiency, particularly between the long and short edges of the pixel. To investigate this observation, we further probed the in-pixel performance by using an attenuation-tunable infrared laser as an accessible proxy for detection efficiency and charge collection efficiency.
@@
Based on the XY-scan results as shown in Fig. 13, four distinct in-pixel positions were selected and marked: (A) the pixel center, (B) the center of the long edge, (C) the center of the short edge, and (D) the diagonal center of $2\times 2$ neighbouring pixels. These four positions, annotated in Fig. 13, were representative of distinct in-pixel locations and were selected to investigate spatial variations in detection efficiency.
@@
As the infrared laser penetrates the silicon bulk and generates electrons and holes along the way, we assessed the pixel’s detection efficiency in response to it to mimic the pathways of charged particles. An RS scan was performed over 100,000 frames (40 seconds) under a 25 Hz laser operation for each reverse bias value across the four in-pixel positions. Given that 1000 laser pulse triggers are expected during this period, the detection efficiency is calculated as the ratio of the measured hit count to 1000, thereby quantifying the pixel’s ability to detect individual laser pulses.
@@
The results shown in Fig. 14 reveal distinct trends across the four in-pixel positions, providing information about the in-pixel variance. Position A (the pixel center) consistently exhibits the highest efficiency, while Position D (the corner) shows the lowest, reflecting diminished charge collection efficiency at the pixel boundary. Position B (long-edge center) and C (short-edge center) display intermediate efficiencies. Notably, the efficiency at Position B is higher than at Position C. This aligns with findings from the JadePix-3 beam test measurement [6], which also revealed a higher detection efficiency at the long-edge center compared to that at the short-edge center. Furthermore, the detection efficiency increases significantly with reverse bias across all positions. Applying a bias of just –2 V is sufficient to raise the efficiency at the center and edge positions (A, B, and C) to nearly 100%, underscoring the performance improvement enabled by the expanding depletion volume. It is also noteworthy that an efficiency drop is observed at Position D when the bias increases from -5 V to -6 V. This may be caused by a smaller lateral depletion region under such a high bias condition, highlighting the importance of carefully selecting the operational bias voltage.
@@
Figure 14: Detection efficiency in response to laser beam as a function of the reverse bias $V_{\rm sub}$. Measured at four in-pixel spots A, B, C, and D as marked in Fig. 13.
@@ 3.6 Optimal Reverse Bias Setting
A comprehensive survey of the JadePix-3’s performance over substrate biases from 0 to –6 V indicates that, under the tuning method presented in this paper, a reverse bias of approximately –5 V offers the best overall trade-off. This optimal working-point not only improves detection efficiency and charge collection ability, but also maintains a low fake-hit rate and input capacitance.
@@ 4 Conclusions and Outlook
These results acquired in laboratory conditions will also provide critical insights to inform subsequent beam test configurations, parameter optimization, as well as the radiation hardness test. Moreover, this work will serve as the architectural and procedural foundation for the tuning and optimization of the next sensor in the JadePix series, thereby driving forward MAPS development to meet CEPC’s requirements.

### v3

@@ Abstract
The Circular Electron-Positron Collider (CEPC), a proposed next-generation $e^{+}e^{-}$ collider to enable high-precision studies of the Higgs boson and potential new physics, imposes rigorous demands on detector technologies, particularly the vertex detector. JadePix-3 is a prototype Monolithic Active Pixel Sensor (MAPS) designed for the CEPC vertex detector. This paper presents a detailed laboratory-based characterization of the JadePix-3 sensor, focusing on the previously under-explored effects of substrate reverse bias voltage on key performance metrics: charge collection efficiency, average cluster size, and hit efficiency of laser. Systematic testing demonstrated that JadePix-3 operates reliably under reverse bias, exhibiting a reduced input capacitance, an expanded depletion region, enhanced charge collection efficiency, and a lower fake-hit rate. These findings confirm the sensor’s potential for high-precision particle tracking and vertexing at the CEPC while offering valuable references for future iterational R&D of the JadePix series.
@@ 1.1 Overview of JadePix-3 design
The JadePix-3 sensor, developed with TowerJazz 180 nm CMOS imaging technology, covers a pixel matrix consisting of 512 rows and 192 columns. The sensor is divided into four distinguishable sectors in the column direction. While the row pitch is uniformly set at 16 $\rm{\upmu m}$ across all four sectors, the column pitch varies between 23.11 $\rm{\upmu m}$ in sector 2 and 26 $\rm{\upmu m}$ in sectors 0, 1, and 3. The variation in the column pitch arises from the implementation of different digital logic blocks within the pixels. These blocks were designed to evaluate the performance of RS latches versus D flipflops, as well as the inclusion or exclusion of a MASK configuration register, aiming to minimize the pixel layout. The functional block diagram of the pixel is illustrated in Fig. 1, while the key differences between the sectors are summarized in Tab. 2.
Figure 1: Pixel Schematic of the JadePix-3 sensor, illustrating the main components: the collection diode, the analog front-end and the digital readout logic. The digital logic includes configuration bits for electrical pulse test (D latch1) and pixel masking (D latch2). The “Hit Register” (blue dashed box) represents the different implementations (RS latch or D flipflop) across the sensor sector designs, as detailed in Tab. 2. Further details on the logic operation can be found in Ref. [5] (Section 2.3).
Table 2: Digital logic blocks for different sectors.
Sector pixel layout digital logic block Remarks
0 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 RS latches
1 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V1 D flipflop, w/o MASK
2 16 $\times$ 23.11 $\rm{\upmu m^{2}}$ DGT_V2 RS latches, w/o MASK
3 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 Same as Sector 0 but with increased current for the analog front-end
@@ 1.2 Characterization Motivation
Figure 3: Circuit topology for the JadePix-3 analog front-end [5]. This analog front-end was adapted from the ALPIDE [7] sensor, featuring low power consumption of 40 nW and minimum threshold below 100 $e^{-}$.
The performance of MAPS is significantly influenced by the bias voltage applied to the substrate. This bias has a multifaceted impact: on one hand, it directly increases the size of the depletion zone and reduces the sensor’s input capacitance [8]. On the other hand, it also affects the characteristics of the integrated circuitry; specifically, as the substrate bias becomes more pronounced, the threshold of NMOS transistors within the analog front-end circuitry (M3, M5, M8 and M9 in Fig. 3) tends to increase due to the body effect, consequently altering the circuit’s operating status. Therefore, adjusting the circuit’s operating point after applying substrate bias and investigating the resulting impact on these key performance metrics becomes essential. These metrics are vital for the R&D of the CEPC vertex detector.
@@ 2.2 Working-point Selection
The operational parameters of the pixels—threshold, fixed pattern noise (FPN), and temporal noise (TN)—were characterized using the standard “S-curve” method. This technique involves injecting a series of electrical pulses of increasing charge ($Q_{\rm injected}$) into each pixel and measuring the hit response. The injected charge is regulated by an external voltage pulse, $V_{\rm pulse}$, according to the relation $Q_{\rm injected}=V_{\rm pulse}\cdot C_{\rm test}$, where $C_{\rm test}$ is the injection capacitor with a designed value of 0.18 fF. The resulting S-shaped curve for each pixel is then analyzed to extract its specific threshold and noise characteristics. The final threshold value is converted to electrons via the value of $C_{\rm test}$.
@@ 0 V 400 500 1307 722
The threshold distributions for these target working-points are presented in Fig. 5. It is observed that these distributions at non-zero reverse biases are noticeably broader compared to zero bias, indicating increased pixel-to-pixel threshold variations. This broadening of the threshold distributions is quantified in Fig. 6, which plots the FPN as a function of bias voltage. The corresponding TN analysis is shown in Fig. 6. An increase in TN is observed under reverse bias conditions. The underlying reasons for this behavior will be discussed in Section 3.2 along with the change of the OUT_A signal amplitude.
@@ 3.2 Input Capacitance
The two methods provide complementary insights. While Method 1 directly captures the amplitude response at a fixed pulse voltage, Method 2, by contrast, keeps the amplitude-dependent $f_{\rm gain}$ factor fixed via maintaining the amplitude value constant, which offers a more reliable $C_{\rm effective}$ result with increasing reverse bias voltage.
Results for these two methods after calculation are shown in Fig. 9. In both methods, the input capacitance can be reduced by more than 50% when going from $V_{\rm sub}=0~{\rm V}$ to $V_{\rm sub}=-2~{\rm V}$, corresponding to the rapid expansion of the depletion region. This reduction is very important as smaller input capacitance increases the charge-voltage conversion coefficient at the input of the front-end circuit for a given charge. As a consequence of the increased charge-voltage conversion, the amplitude of the analog signal (OUT_A) increases significantly with reverse bias. In Section 2.2.1, to keep the input-referred threshold constant, the VCASN was tuned at different reverse bias voltage. The tuning of VCASN actually lowers the effective OUT_A baseline which is relative to the voltage required to switch the transistor M9 on. The increase of FPN and TN in Fig. 6 and 6, respectively, may be related to the tuning of VCASN because the same effect observed in a dedicated VCASN scan shown in Ref. [5] (Fig. 13).
@@ 3.3 Fake-hit Rate Test
JadePix-3 employs both rolling shutter (RS) and global shutter (GS) readout sequence modes. In the conventional rolling shutter mode, one row is activated for reading at a time. This process requires 192 ns to transmit hit data and reset the pixels, and scanning the entire 512-row matrix takes 98.3 $\rm{\upmu s}$. While in global shutter, all 512 rows are frozen for readout after a period of acquisition time. To compare the two readout sequence modes, the acquisition time of the GS mode is set equivalent to 98.3 $\rm{\upmu s}$ and the fake-hit rate for each sector $i$ in both modes can be calculated using a unified formula:
@@
where $N_{i}^{\text{mode}}$ represents the total number of false hits in sector $i$, $N_{\text{frames}}^{\text{mode}}$ is the number of frames collected, and $N_{\text{pixels}}=512\times 48$ is the number of pixels per sector. For RS mode, the test system can readout continuously while for the GS mode the test system has to launch one readout for one frame, which limits the total number of frames that can be collected in the test. In the absence of external sources, a total of 100,000 frames ($N_{\text{frames}}^{\text{RS}}=100,000$) are collected in the RS mode, while 1,000 frames ($N_{\text{frames}}^{\text{GS}}=1,000$) are collected in the GS mode.
The sensitivity limit is defined as the lowest measurable fake-hit rate in the measurement and is determined by the number of frames for these two modes. Based on Eq. 3, this results in a sensitivity limit of $\rm 4\times 10^{-10}$ for the RS mode and $\rm 4\times 10^{-8}$ for the GS mode.
@@ 3.4 Radioactive Source Test
Two radioactive sources, $^{90}\rm Sr$ and $^{55}\rm Fe$, were employed to assess the sensor’s response to external ionizing radiation. $^{90}\rm Sr$, a $\beta$ emitter, generates electrons that traverse the sensor and deposit charge over an extended region. This distributed ionization is particularly sensitive to changes in the sensor’s depletion characteristics and charge-sharing behavior under reverse bias, thereby providing insights into the sensor’s charge collection efficiency. $^{55}\rm Fe$ emits 5.9 keV X-rays that interact with silicon primarily via the photoelectric effect and deposit an ionization charge of 1640 $e^{-}$ in a small point-like spatial region compared to the dimensions of the sensor, resulting in highly localized charge deposition. This characteristic makes $^{55}\rm Fe$ an ideal probe for evaluating the sensor’s response to concentrated ionization, providing a signal whose magnitude is comparable to that of a minimum ionizing particle (about 66 electron-hole pairs are generated per micron [11] in the 20 $\rm{\upmu m}$ epitaxial layer of the JadePix-3 prototype).
@@ 3.5 Infrared Laser Test
Following the configuration of the laser test platform outlined in Section 2.1, a systematic investigation of the in-pixel laser hit efficiency was performed using an infrared laser beam incident from the backside of the chip. The laser operates at a wavelength of 1064 nm. The laser’s pulse energy is adjustable between 0 and 60 pJ, modulated via an attenuation factor ranging from 0% tune (maximum pulse energy) to 100% tune (minimum pulse energy), controlled through the Laser Power Tune function of the system. For this laser measurement, the selected pixel is located in sector 2, which has the smallest pixel pitch among the sectors (see Tab. 2).
@@ 3.5.2 In-Pixel Hit Efficiency
Previous beam test results [6] indicated a non-uniform in-pixel detection efficiency, particularly between the long and short edges of the pixel. To investigate this observation, we further probed the in-pixel performance by using an attenuation-tunable infrared laser as an accessible proxy for hit efficiency and charge collection efficiency.
@@
Based on the XY-scan results shown in Fig. 14, four distinct in-pixel positions were selected and marked: (A) the pixel center, (B) the center of the long edge, (C) the center of the short edge, and (D) the diagonal center of $2\times 2$ neighbouring pixels. These four positions, annotated in Fig. 14, were representative of distinct in-pixel locations and were selected to investigate spatial variations in hit efficiency.
@@
As the infrared laser penetrates the silicon bulk and generates electrons and holes along the way, we assessed the pixel’s hit efficiency in response to it to mimic the pathways of charged particles. For this laboratory measurement, we define the hit efficiency as the probability that an incident laser pulse results in a registered hit above the threshold. The laser intensity was kept the same as in the preceding XY-scan, but the pulse frequency was lowered to 25 Hz to ensure that a maximum of one laser pulse could arrive within a single frame. A 40-second Rolling Shutter scan was performed for each reverse bias value across the four in-pixel positions. Given that 1000 laser pulse triggers are expected during this period, the hit efficiency is calculated as the ratio of the measured hit count to 1000, thereby quantifying the pixel’s ability to detect individual laser pulses.
@@
The results shown in Fig. 15 reveal distinct trends across the four in-pixel positions, providing information about the in-pixel variance. Position A (the pixel center) consistently exhibits the highest efficiency, while Position D (the corner) shows the lowest, reflecting diminished charge collection efficiency at the pixel boundary. Position B (long-edge center) and C (short-edge center) display intermediate efficiencies. Notably, the efficiency at Position B is higher than at Position C. This aligns with findings from the JadePix-3 beam test measurement [6], which also revealed a higher hit efficiency at the long-edge center compared to that at the short-edge center. Furthermore, the hit efficiency increases significantly with reverse bias across all positions. Applying a bias of just –2 V is sufficient to raise the efficiency at the center and edge positions (A, B, and C) to nearly 100%, underscoring the performance improvement enabled by the expanding depletion volume. It is also noteworthy that an efficiency drop is observed at Position D when the bias increases from -5 V to -6 V. This may be caused by a smaller lateral depletion region under such a high bias condition, highlighting the importance of carefully selecting the operational bias voltage.
@@
Figure 15: Hit efficiency in response to laser beam as a function of the reverse bias $V_{\rm sub}$. Measured at four in-pixel spots A, B, C, and D as marked in Fig. 14. The laser intensity was the same as that used for the XY-scan, but the frequency was reduced to 25 Hz to ensure at most one pulse per frame.
@@ 3.6 Optimal Reverse Bias Setting
A comprehensive survey of the JadePix-3’s performance over substrate biases from 0 to –6 V indicates that, under the tuning method presented in this paper, a reverse bias of approximately –5 V offers the best overall trade-off. This optimal working-point not only improves hit efficiency and charge collection ability, but also maintains a low fake-hit rate and input capacitance.
@@ 4 Conclusions and Outlook
These results acquired in laboratory conditions will also provide critical insights to inform subsequent beam test configurations, parameter optimization, as well as the radiation hardness test. Moreover, this work will serve as the architectural and procedural foundation for the tuning and optimization of the next sensor in the JadePix series, thereby driving forward MAPS development to meet CEPC’s requirements, including the ultimate 3 $\rm{\upmu m}$ spatial resolution target. Achieving this will require further dedicated R&D, commencing with extensive beam tests to quantify the impact of the reverse-bias settings explored in this study on the spatial resolution. A critical part of this future work will be to pinpoint the root causes of degradation in position resolution under different reverse bias voltages—whether stemming from the chip design, experimental setup, or analysis algorithms—in order to guide the next sensor iteration.